For each of the problems below solve for monopoly quantity, monopoly price, total revenue, total cost, profit, and dead weight loss.

a. Qd = 100 – 0.5p                           TC = 2Q2

b. Qd = 500 – 0.2p TC = Q2 + 100Q

c. Qd = 30 – 0.6p                             TC = Q2 + 10Q +100

d. Qd = 1000 – 0.25p                       TC = 2Q2 + 25Q + 100
Answer #1

a Qd = 100- 0.5P Now, invu se demand 100-Qd = 0.5 P . =) 200 - 2 Qd =P Total revenue (TR) PQd = (200-2Qd)? = 200 Qd = 2 Qd² -Profit T = Total revenue - Total cost = $ (3750 - 1250) - $ 2500 Now, p= 200-2 ad When po, Qd = 100 When Qdzo p= 200 MarginalTo get point ce we equation 2 P= MC MC = 4Qd we know Qd = 25 So Mc- 25×4 = 6100 Now P= mc boo implin 200-2 Qd = 100 100 = 2Qb) Qd = 500-012P TC=Q² + 1000Q TC=Q² + 100 Q b Qd = 500- 0.2P =) Now, ad = 500-0.2P = 500- Qd = 0.2P 2500 - 5Qd = pi Total reTo tal cost of an To - Q²+ 100Q - (200) 27100(200) :-$60,000 Profit will be a TR- TC) =47300,000 - 60,000) $ 240, oor. I Now,Puce mo sa MC =2Qtlo 2500 1500 500 100 MR 200 250 S . Фuox8* is the area of A ABC Scanned with loss Dead brught.Now, to get point B we substitute Qd=200 in the MR function Tus, MR = 2500-10Qd ) = 2500 – 10 (200) 2 500 Now. Deadweight lTe= Q 7 1004100 C) Qd = 30-0.6P Now Qd = 30-0.6 P. =) 30-Q d = 0.6P =) 50 - 1.66 Qd = p. Now Total revenue . TR = P Qd = 50 QTotul cost Qud + load +100 = 1715273+1017152)+100 = $ 231.75 profit will be = TR-TC I - $ (2 82.15 - 232.15) = $50 Now Qd = 3Now to get Substitute equation point Qd- & we 7.52 in MR They MR = 50-3.32 Qd = so- 3.32 (7152) = 25 Thes, Deadweight loss...Equating MR = me we get 0675 l 4000-8Qd = 4Q tas - - 12 Qd - 3975 =) Qd = 331.25 Now p= 4000 - 4Qd. - 4000 - 41351.95) = $267Now I Qd = 1000 - 0.25 P When Qd = 0, P= $4000 when p=0 => Qd = 1000 osti MR= 4000 - 8 Qd When Qd =0, MR 4000 When MR = 0 QdNow, Deadweight loss - Area of A ABC - { x ( 500- 33).25) X (2675-1350) = { x 168.95 x 13.85 = $ 111,929.38 Sanped with